calculate the osmotic pressure of a solution containing 15.0 grams of Ca(NO3)2 (molar mass=164.088g/mol) in 250.ml of solution at 25 degrees C assume complete dissociation Ca(NO3)2
1
Expert's answer
2019-12-09T08:05:27-0500

 Osmotic pressure, p=?

Mass m=15g

Molar mass, M=164.088g/mol.

Volume, v=0.00025m3.

Temperature, T=298.15K.

"P=(nRT)\/v"

Number of moles "n=m\/M"

                                n= 15/164.088 =0.09141moles

p= (0.09141*8.314*298.15)/0.00025.

p= 906355.42pa